Grade 12th passMechanics

If g is the inverse of f and f'(x)=1/(1+x^3). then g'(x) is equal to?

Ajay
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
 
g is inverse of function f, then
f(g(x)) = x
differentiating wrt x
f’(g(x))*g’(x) = 1
So g’(x) = 1/f’(g(x)) = 1 + {g(x)}3
Hope it clears. If you like answer then please approve it
 
 
Regards
Arun (askIITians forum expert)